OpenStudy (anonymous):

Input the equation for the following line. Reduce all fractional answers to lowest terms. A line that passes midway between and is parallel to both x - y = 6 and x - y = -6.

OpenStudy (anonymous):

y = x ?

OpenStudy (anonymous):

I agree

OpenStudy (anonymous):

no it says that wrong

OpenStudy (anonymous):

we are looking for a line of the form y=mx+b, where m is the slope and b is the y intercepts

OpenStudy (anonymous):

ok i'll graph it

OpenStudy (anonymous):

you can rearrange the given lines so that they are of the same form

OpenStudy (anonymous):

yep - its correct

OpenStudy (anonymous):

one passes thru y = 6 and other thru y = -6 0 is at the center

OpenStudy (anonymous):

the two lines are y=x+6 and y=x-6, so we know the slope of the line we are seeking must be the same as these two because it is given that they are parrallel. (parallel means the same slope).

OpenStudy (anonymous):

so the slope of te line we seek is 1

OpenStudy (anonymous):

right

OpenStudy (anonymous):

i'm afraid your given answer is wrong mih

OpenStudy (anonymous):

the two lines we are given are both 6 units away from y=0 at x=0 so that means that at x=0 the midpoint between the lines is (0,0)

OpenStudy (anonymous):

so our line is y=1x+0 or simply y=x

OpenStudy (anonymous):

yep